if the altitude of a triangle is doubled and the base is multiplied by 3, what is the effect on the area?
Answer by drk(1908)

The formula for triangle area is
A = (1/2)bh
we change this based on given information to
A=(1/2)(3b)(2h)
a= (1/2)(bh)(6)
This means the new triangle has an area 6 times bigger.
--
example:
original: b = 2, h = 5
A = (1/2)(2*5) = 5
new: b = 6, h = 10
A = (1/2)(6*10) = 30
New area = 6 times old area.